Assume that the demand for capital is determined by the equation K* = 0.25 Y /rc. The

nominal interest rate is %12, the inflation rate is %6 percent, and

capital depreciates at %10 percent per year.

a. Initially, income is $16,000. What is the desired capital stock?

b. Now income doubles. What is the new desired capital stock?

c. Given the simple accelerator model of investment, what is the rate of net investment?

Gross investment?


1
Expert's answer
2021-06-10T12:57:19-0400

a) rc = rental cost of capital = r + d


r = real interest rate

d = depreciation rate


As According to fisher equation, i = π + r

π : inflation rate & i = nominal interest rate

So r = i- π = 12-6 = 6%

"so rc = 6\% + 10\% = 16\%"


thus K "= \frac {(0.25×16,000)} {0.16}



= 25,000"


b) now Y = 32,000


K = "\frac {(0.25×32,000)} {0.16}



= 50,000"


C) NEW rc = 32%


so new K = "\frac {(0.25×16,000)} {0.32}



= 12,500"